OpenStudy (anonymous):

A student is running a 3-kilometer race. He runs 1 kilometer every 2 minutes. Select the function that describes his distance from the finish line.

OpenStudy (anonymous):

f(x) = - 1 over 3x + 2 f(x) = - 1 over 2x + 3 f(x) = 1 over 2x + 3 f(x) = 1 over 3x + 2